OpenStudy (anonymous):

Can someone help me? Im stuck on an expression problem. It asks to simplify 5/8 Y - 3/8 Y + 2Y

OpenStudy (zzr0ck3r):

can you tell me \[\frac{5}{8} - \frac{3}{8} + 2 = ?\]

OpenStudy (zzr0ck3r):

first you want a common denominator which is 8 so \[2=\frac{?}{8}\]

OpenStudy (anonymous):

Factor out the Y and then add up the coefficients.

OpenStudy (zzr0ck3r):

im not sure if they know how to factor at the level.